OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: 2x + 2y = 10 x = –y + 5 A. There are an infinite number of solutions. B. There is no solution. C. The solution is (4, –1). D. The solution is (2, –3). im thinking its A but just to check is it?

OpenStudy (anonymous):

x = -y + 5 so sub in -y + 5 in for x in the other equation 2x + 2y = 10 2(-y + 5) + 2y = 10 --- distribute -2y + 10 + 2y = 10 -- subtract 10 from both sides -2y + 2y = 10 - 10 0 = 0 when your answer comes out equal, there is infinite solutions. If your answer would have come out like this : 2 = 3, then there would have been no solution. INFINITE SOLUTIONS is the answer

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

no problem :)